OpenStudy (anonymous):

Which equation would represent the line with a slope of zero and the point (2,4)? x = 2 x = 4 y = 2 y = 4

OpenStudy (muhammad_nauman_umair):

y=4

OpenStudy (muhammad_nauman_umair):

As y-y1=m(x-x1) here m=0 so y-y1=0 y=y1 y=4